OpenStudy (anonymous):

P4O10 + 6PCl5---> 10 POCl3 How many grams of POCl3 are produced when 225.0g of P4O10 and 675.0g of PCl5 react? How would I find this using dimensional analysis!

OpenStudy (aaronq):

first you need to identify the limiting reactant. Do this by converting the masses given to moles and dividing by their stoichiometric coefficients. Whichever has less moles after the division, is the limiting reactant. Use the amount of moles of the limiting reactant to find the moles of product made. Use the stoichiometric coefficients to build a ratio, like this: e.g. for a general reaction: \(\color{red}{a}A + \color{blue}{b}B\) \(\rightleftharpoons\) \( \color{green}{c}C\) where upper case are the species (A,B,C), and lower case (a,b,c) are the coefficients, \(\dfrac{n_A}{\color{red}{a}}=\dfrac{n_B}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\) To convert mass to moles, use the relationship: \(n=\dfrac{m}{M}\) where, M=molar mass, m=mass, and n= moles.